Data for Adelphia Inc. for May 2018 is provided below.

Adelphia expected to produce and sell 11,000 units of product, each requiring 0.67 pounds of aluminum at a cost of $30.41 per pound. Adelphia produced and sold 11,400 units of product and used 8,200 pounds at a cost of $248,700.

Adelphia expected that each unit of product would require 1.2 direct labor hours at a cost of $10.25 per hour. Actual direct labor cost was $10.59 per hour, for a total of $142,420.

In May Adelphia incurred $17,300 of variable manufacturing. Overhead was applied based on direct labor hours at a rate of $1.34 per hour.

Required:

Calculate the following variances and indicate for each, whether it is favorable or unfavorable: Material price variance

Labor rate variance

Material quantity variance

Labor efficiency variance

Material cost (total) variance

Labor cost (total) variance

Total variable overhead (controllable) variance
Answer #1
Particulars Computations
1 Material Price Variance = Actual cost of purchase of material - Budgeted Cost of purchase of material
= $248700- $(8200*30.41)
= $248700-$249362
=$662 (Favorable)
2 Labor Rate Variance = Actual Amt paid to labors - Standard Cost of Actual Hours
= $142420- $137847(*working note-1)
=$4572 (Unfavorable)
3 Material Quantity Variance (Actual usage of material - Standard usage)*Standard cost per unit
=(8200-7638**(WN-2))*30.41
=$17090 (Unfavorable)
4 Labor efficiency variance = Standard Cost of Actual Hours - Standard Cost
= $137847 - $140220(***WN-3)
=$ 2373 (favorable)
5 Material cost (total) variance = Material price variance + Material quantity variance
= 662+(-17090)
= $16428 (Unfavorable)
6 Labor cost (total) variance =Labor Rate Variance+ Labor efficiency variance
= -4572+2373
=$ 2199 (Unfavorable)

*Working note(WN)-1

Actual hours worked=$142420/10.59 = 13448 hours

Standard Cost of Actual Hours= expected per hour cost*actual hours worked = 13448*10.25 =$137847

**Working note(WN)-2

Standard usage=Actual unit used*budgeted pounds per unit = 11400*0.67 = 7638 pounds

**Working note(WN)-3

Standard hours= Actual units*expected hours per unit = 11400*1.2 = $13680

Standard cost= standard hours*standard rate = 13680*10.25 = $140220

  

(Kindly comment in case of any further clarification in answer)